Atsie (atsie):

How am I to solve for h? I'm really bad at this, and confused besides. (See pic's below)

Atsie (atsie):



Atsie (atsie):



jhonyy9 (jhonyy9):

g(195-h)=1/2 *v^2 g=32ft/s^2 v= 49 ft/s^2 divide both sides by g and will get 195-h=v^2/(2g) add to both sides h and will get 195 = v^2/(2g) +h subtract from both sides v^2/(2g) and will get 195 - v^2/(2g) = h hope this will help you and will be understandably sure right easy